# Causal Inference: A Tale of Three Frameworks

### v1

@@ Acknowledgements
We are grateful to the Editor for encouraging the development of this work, and to Forrest Crawford for insightful discussions that helped shape several key aspects. We acknowledge the use of OpenAI’s ChatGPT for proofreading and refining the writing.
@@
For binary variables, it is well known that a $2\times 2$ table with fixed marginals has exactly one degree of freedom. Any strictly positive such table can therefore be uniquely represented by a single scalar measuring the strength of association. A standard choice for this scalar is the conditional odds ratio (Chen,, 2007; Tchetgen Tchetgen et al.,, 2018):

### v2

@@ Acknowledgements
We are grateful to the Editor for encouraging the development of this work, and to Forrest Crawford and Sander Greenland for insightful discussions that helped shape several key aspects. We acknowledge the use of OpenAI’s ChatGPT for proofreading and refining the writing.
@@
For binary variables, it is well known that a $2\times 2$ table with fixed marginals has exactly one degree of freedom. Any strictly positive such table can therefore be uniquely represented by a single scalar measuring the strength of association. A standard choice for this scalar is the conditional odds ratio (Chen, 2007; Tchetgen Tchetgen et al., 2018):